Ehrmann Data Systems is considering a project that has the following cash flow and WACC data. What is the project's MIRR? Note that a project's projected MIRR can be less than the WACC (and even negative), in which case it will be rejected.


WACC: 9.00% 

Year 0 1 2 3

Cash flows -$1,000 $450 S450 $450 


14.53% 17.29% 13.70% 13.84% 13.28% 
Answer #1

Answer is 13.84%

WACC = 9.00%

Present Value of Cash Outflow = $1,000

Future Value of Cash Inflows = $450*1.09^2 + $450*1.09 + $450
Future Value of Cash Inflows = $1,475.145

MIRR = (Future Value of Cash Inflows / Present Value of Cash Outflow)^(1/n) - 1
MIRR = ($1,475.145 / $1,000)^(1/3) - 1
MIRR = 1.475145^(1/3) - 1
MIRR = 1.1384 - 1
MIRR = 0.1384 or 13.84%
Answer #2

SOLUTION ;




Period of project, n = 3 years

Initial investment = 1000 ($) (it is cash outflow)

Cash inflows = $450 each year for 3 years.

r = WACC = 9% = 0.09

=> 1 + r = 1.09


FV of positive cash inflows = 450(1.09^2 + 1.09^1 + 1) = 1475.15 ($)

PV of cash outflow = 1000 ($)


So,


MIRR 

= (FV of cash inflows / PV of cash outflows)^(1/n) - 1

= (1475.15/1000)^(1/3) - 1

= 0.1384

= 13.84% .


MIRR is = 13.84%  : 4th Option  (ANSWER).



answered by: Tulsiram Garg